OpenStudy (anonymous):

find the missing side of the right triangle below: [on the left side going up is 10 on the bottom is 24.

OpenStudy (anonymous):

if the third side is the hypotenuse, then H² = P² + B² = (10)² + (24)² = 100 + 576 = 676 H² = 676 H = √676 H = 26

OpenStudy (anonymous):

a^2 + b^2 = c^2 would be the equation you would use. 10^2 + 24^2 = c^2, which would give you 100 + 576 = 676. The square root of 676 is 26, which is your answer.

OpenStudy (anonymous):

thx